# Neural Spacetimes for DAG Representation Learning

### v1

@@ 3.3 Computational Implementation
Representing graph edge directionality as a partial order in the embedding manifold. Given a weighted directed graph $G_{D}=(E_{D},V,W_{D})$, let $x_{u},x_{v}\in\mathbb{R}^{N}$ be the feature vectors associated with the nodes $u,v\in V$, which we want to embed as events in our spacetime: $\mathcal{S}$ takes as input these features and maps them to an embedding in the manifold. The edge set $E_{D}$ induces a binary non-symmetric adjacency matrix $\mathbf{A}$. In our case, $G_{D}$ is not any digraph but a DAG, hence if $A_{uv}=1\Rightarrow A_{vu}=0$. Both entries can only be equal when they are 0, $A_{uv}=A_{vu}=0$. In the latter case, the nodes may either be causally disconnected, or causally connected but not neighbours. If the input graph does not satisfy these properties our spacetime embedding parametrized by the NST will not be able to faithfully embed the input graph in terms of its edge directionality.
@@
Representing graph edge weights as distances in the embedding manifold. Likewise, $W_{D}$ induces a distance matrix $\mathbf{D}$ with entries $D_{uv}$ being the causal distance between events $u,v$. Although theoretically this could be computed as the shortest path graph geodesic distance for two causality connected and distant $u$ and $v$ (similar to equation 1 but for weighted directed graphs, see Appendix D.1), in practice we only optimize for the one-hop neighbourhood. In particular, if $u=v\Rightarrow D_{uv}=0$, if $u\preccurlyeq v\land v\in\mathcal{N}(u)\land u\neq v\Rightarrow D_{uv}>0$. The distance $D_{uv}$ is ignored otherwise, that is, we do not model the distance between nodes in the original graph that are causally connected but outside the one-hop neighbourhood of each other nor do we model the distance between events that are not causality connected (anti-chains). This is in line with the literature on graph construction via Lorentzian pre-length spaces (Law & Lucas, 2023). In the case of NSTs, it is achieved using the connectivity $A_{uv}$ as a mask in the loss function.
@@
Local geometry optimization and global geometry implications. Note that although we optimize for the one-hop neighbourhood of each node only, transivity of the causal connectivity of nodes across hops will be satisfied by definition of the partial order (Remark 3). Additionally, although the partial order between anti-chains is not directly optimized, as the number of time dimensions increases, it is increasingly probable that $\lesssim^{\mathcal{T}}$ will not be satisfied for nodes not causally connected, as desired. Conversely, there is no guarantee that when directly evaluating the distance between causally connected nodes using $\mathcal{D}$ we will obtain the graph geodesic distance between nodes. In summary, if $x_{u},x_{v},x_{w}\in\mathbb{R}^{N}$ are feature vectors for nodes $u,v,w\in V$, and $\hat{x}_{u},\hat{x}_{v},\hat{x}_{w}\in\mathbb{R}^{D+T}$ are their $\mathcal{E}$ encodings in the intermediate Euclidean space used by the NST, if $x\lesssim^{\mathcal{T}}y\land y\lesssim^{\mathcal{T}}z\Rightarrow x\lesssim^{\mathcal{T}}z$. On the other hand, in general $\mathcal{D}(x,y)+\mathcal{D}(y,z)=D_{uv}+D_{vw}$, and provided that this particular path corresponds to the graph geodesic $D_{uv}+D_{vw}=d_{G}(u,w)$, but $\mathcal{D}(x,y)+\mathcal{D}(y,z)\neq\mathcal{D}(x,z)$. The latter inequality is acceptable and not required to faithfully encode the graph edge weights.
@@
Training and Loss Function. Let us use the matrix $\mathbf{X}\in\mathbb{R}^{M\times N}$ to denote the collection of feature vectors for $M=|\mathcal{V}|$ nodes associated with the DAG, $G_{D}$. The neural spacetime parameters are optimized with respect to the following loss:
$\mathcal{L}_{G_{D}}(\mathbf{X},\mathbf{A},\mathbf{D})\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\sum_{u=u_{1}}^{u_{M}}\sum_{v=v_{1}}^{v_{M}}\mathcal{L}_{uv}\left(\mathcal{S}(x_{u},x_{v}),(A_{uv},D_{uv})\right),$ (6)
@@
where $x_{u}$ extracts the row feature vector for $u$ from matrix $\mathbf{X}$. The goal is to learn appropriate embeddings so that nodes connected by directed edges are represented as causally connected events in the time dimensions of the manifold, and respect the distance $D_{uv}$ induced by the graph weights in the space dimensions. To do so, we can further divide the loss function into a (causal) distance loss, $\mathcal{L}^{D}_{uv}$, and a causality loss, $\mathcal{L}^{C}_{uv}$. We remind the reader of the definition $\hat{x}_{u},\hat{x}_{v}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\mathcal{E}(x_{u}),\mathcal{E}(x_{v})\in\mathbb{R}^{D+T}$. Hence the loss can be partitioned into:
$\mathcal{L}_{uv}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\mathcal{L}^{D}_{uv}\left(\mathcal{D}(\hat{x}_{u},\hat{x}_{v}),(A_{uv},D_{uv})\right)+\mathcal{L}^{C}_{uv}\left(\mathcal{T}(\hat{x}_{u}),\mathcal{T}(\hat{x}_{v}),A_{uv}\right).$ (7)
@@
Next, we dissect each of the losses. The distance loss for a pair of nodes or events is a standard mean squared error (MSE) loss between the predicted and ground truth distance, with masking given by the adjacency matrix:
@@
$\smash{\mathcal{L}^{D}_{uv}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}A_{uv}\textrm{MSE}\left(\mathcal{D}(\hat{x}_{u},\hat{x}_{v}),D_{uv}\right).}$ (8)
On the other hand, the causality loss (with respect to the one-hop neighbourhood) is
$\mathcal{L}^{C}_{uv}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\,A_{uv}\mathcal{L}^{*}_{C}\Big{(}\sum_{t=1}^{T}\,\operatorname{SteepSigmoid}(\mathcal{T}(\hat{x}_{u})_{t}-\mathcal{T}(\hat{x}_{v})_{t})\Big{)},$ (9)
@@
where $\mathcal{L}^{*}_{C}$ is a function that takes as input the expression above and $\operatorname{SteepSigmoid}(x)=\frac{1}{1+e^{-10x}}$ (we omit some details here for simplicity, see Appendix C.3 for a lengthier explanation). The loss for two causally connected events $u\preccurlyeq v$ in the first neighbourhood of each other ($A_{uv}=1$) is minimized when $\hat{x}_{u}\lesssim^{\mathcal{T}}\hat{x}_{v}$ is satisfied (equation 5). If the time coordinates of $\hat{x}_{v}$ associated with the event $v$ are elementwise greater than all the time coordinates of $\hat{x}_{u}$ for $u$ then all $\operatorname{SteepSigmoid}$ activation functions will return $\approx 0$.
4 Experimental Validation
In this section, we provide an overview of our experimental procedure and results.
@@ 4 Experimental Validation
Synthetic Weighted DAG Embedding. We generate DAGs (see Appendix D.2) embedded in the 2D plane with associated local distances between nodes given by several metrics. We measure the embedding capabilities of NSTs compared to closed-form spacetime geometries such as the Minkowski and De Sitter spaces in Law & Lucas (2023). Note that although we do not optimize the geometries in the case of the baselines, we do use a neural network encoder to map points to events in the manifolds. We quantify both the average and maximum metric distortion (the ratio between true and predicted distances) following Kratsios et al. (2023b), as well as the accuracy of the time embedding. As we can see in Table 1, we are always able to embed edge directionality (0 for no edges embedded correctly, 1 for all edges embedded correctly). In terms of metric distortion, we can see that as the embedding dimension increases, both average and maximum distortion decrease. NSTs are particularly good at retraining low distortions in low-dimensional embedding spaces.
@@
Real-World Network Embedding. Next, we test our approach on real-world networks. In Table 2, we present results for the Cornell, Texas, and Wisconsin (WebKB) datasets (Rozemberczki et al., 2021), which are based on web pages represented as nodes and directed hyperlinks between them. In this case, all nodes have bag-of-word features, which are used as input to the neural spacetime encoder. The neural partial order encodes the hyperlink directionality between websites, and the neural (quasi-)metric learns the connectivity strength as the cosine similarity between connected web page features. We achieve very low distortions, showcasing the embedding capabilities of our network. Note that, from a metric learning perspective, real-world datasets are generally less challenging than the metrics presented in Table 1, which were chosen to be particularly unconventional on purpose. In terms of the time embedding, we manage to mostly encode directionality; however, these datasets are not pure DAGs, so it is not possible to embed them perfectly. We also work with real-world gene regulatory network datasets (Marbach et al., 2012) in line with previous spacetime representation learning literature (Law & Lucas, 2023; Sim et al., 2021). We achieve both good spatial and causal embeddings. For further details refer to Appendix D.3.
Experimental details and hyperparameters for all setups can be found in Appendix D. Additionally, we present tree (spatial only) embedding experiments comparing neural snowflakes to neural (quasi-)metrics in NSTs, as well as hyperbolic neural networks (HNNs), in Appendix D.1.

### v2

@@ 3.2 Computational Implementation
Representing graph edge directionality as a partial order in the embedding manifold. Given a weighted directed graph $G_{D}=(E_{D},V,W_{D})$, let $x_{u},x_{v}\in\mathbb{R}^{N}$ be the feature vectors associated with the nodes $u,v\in V$, which we want to embed as events in our spacetime: $\mathcal{S}$ takes as input these features and maps them to an embedding in the manifold. The edge set $E_{D}$ induces a binary non-symmetric adjacency matrix $\mathbf{A}$. In our case, $G_{D}$ is not any digraph but a DAG, hence if $A_{uv}=1\Rightarrow A_{vu}=0$. Both entries can only be equal when they are 0, $A_{uv}=A_{vu}=0$. In the latter case, the nodes may either be causally disconnected, or causally connected but not neighbors. If the input graph does not satisfy these properties our spacetime embedding parametrized by the NST will not be able to faithfully embed the input graph in terms of its edge directionality.
@@
Representing graph edge weights as distances in the embedding manifold. Likewise, $W_{D}$ induces a distance matrix $\mathbf{D}$ with entries $D_{uv}$ being the causal distance between events $u,v$. Although theoretically this could be computed as the shortest path graph geodesic distance for two causality connected and distant $u$ and $v$ (similar to equation 1 but for weighted directed graphs, see Appendix D.1), in practice we only optimize for the one-hop neighborhood. In particular, if $u=v\Rightarrow D_{uv}=0$, if $u\preccurlyeq v\land v\in\mathcal{N}(u)\land u\neq v\Rightarrow D_{uv}>0$. The distance $D_{uv}$ is ignored otherwise, that is, we do not model the distance between nodes in the original graph that are causally connected but outside the one-hop neighborhood of each other nor do we model the distance between events that are not causality connected (anti-chains). This is in line with the literature on graph construction via Lorentzian pre-length spaces (Law & Lucas, 2023). In the case of NSTs, it is achieved using the connectivity $A_{uv}$ as a mask in the loss function.
@@
Local geometry optimization and global geometry implications. Note that although we optimize for the one-hop neighborhood of each node only, transivity of the causal connectivity of nodes across hops will be satisfied by definition of the partial order (Remark 4). Additionally, although the partial order between anti-chains is not directly optimized, as the number of time dimensions increases, it is increasingly probable that $\lesssim^{\mathcal{T}}$ will not be satisfied for nodes not causally connected, as desired. Conversely, there is no guarantee that when directly evaluating the distance between causally connected nodes using $\mathcal{D}$ we will obtain the graph geodesic distance between nodes. In summary, if $x_{u},x_{v},x_{w}\in\mathbb{R}^{N}$ are feature vectors for nodes $u,v,w\in V$, and $\hat{x}_{u},\hat{x}_{v},\hat{x}_{w}\in\mathbb{R}^{D+T}$ are their $\mathcal{E}$ encodings in the intermediate Euclidean space used by the NST, if $(\hat{x}_{u}\lesssim^{\mathcal{T}}\hat{x}_{v})\land(\hat{x}_{v}\lesssim^{\mathcal{T}}\hat{x}_{w})\Rightarrow\hat{x}_{u}\lesssim^{\mathcal{T}}\hat{x}_{w}$. On the other hand, in general $\mathcal{D}(\hat{x}_{u},\hat{x}_{v})+\mathcal{D}(\hat{x}_{v},\hat{x}_{w})=D_{uv}+D_{vw}$, and provided that this particular path corresponds to the graph geodesic $D_{uv}+D_{vw}=d_{G}(u,w)$, but $\mathcal{D}(\hat{x}_{u},\hat{x}_{v})+\mathcal{D}(\hat{x}_{v},\hat{x}_{w})\neq\mathcal{D}(\hat{x}_{u},\hat{x}_{w})$. The latter inequality is acceptable and not required to faithfully encode the graph edge weights.
@@
Training and Loss Function. Let us use the matrix $\mathbf{X}\in\mathbb{R}^{M\times N}$ to denote the collection of feature vectors for $M=|\mathcal{V}|$ nodes associated with the DAG, $G_{D}$. The NST minimizes the following loss:
$\mathcal{L}_{G_{D}}(\mathbf{X},\mathbf{A},\mathbf{D})\stackrel{{\scriptstyle\mbox{def.}}}{{=}}\sum_{u=u_{1}}^{u_{M}}\sum_{v=v_{1}}^{v_{M}}\mathcal{L}_{uv}\left(\mathcal{S}(x_{u},x_{v}),(A_{uv},D_{uv})\right),$ (6)
@@
where $x_{u}$ extracts the row feature vector for $u$ from matrix $\mathbf{X}$. The goal is to learn appropriate embeddings so that nodes connected by directed edges are represented as causally connected events in the time dimensions of the manifold, and respect the distance $D_{uv}$ induced by the graph weights in the space dimensions. To do so, we can further divide the loss function into a (causal) distance loss, $\mathcal{L}^{D}_{uv}$, and a causality loss, $\mathcal{L}^{C}_{uv}$. We remind the reader of the definition $\hat{x}_{u},\hat{x}_{v}\stackrel{{\scriptstyle\mbox{def.}}}{{=}}\mathcal{E}(x_{u}),\mathcal{E}(x_{v})\in\mathbb{R}^{D+T}$. Hence the loss can be partitioned into:
$\smash{\mathcal{L}_{uv}\stackrel{{\scriptstyle\mbox{def.}}}{{=}}\mathcal{L}^{D}_{uv}\left(\mathcal{D}(\hat{x}_{u},\hat{x}_{v}),(A_{uv},D_{uv})\right)+\mathcal{L}^{C}_{uv}\left(\mathcal{T}(\hat{x}_{u}),\mathcal{T}(\hat{x}_{v}),A_{uv}\right).}$ (7)
@@
We dissect the two loss terms. The distance loss for a pair of nodes is the mean squared error (MSE) loss between the predicted and ground truth distance, with masking given by the adjacency matrix:
@@
$\smash{\mathcal{L}^{D}_{uv}\stackrel{{\scriptstyle\mbox{def.}}}{{=}}A_{uv}\textrm{MSE}\left(\mathcal{D}(\hat{x}_{u},\hat{x}_{v}),D_{uv}\right).}$ (8)
On the other hand, the causality loss (with respect to the one-hop neighborhood) is
$\mathcal{L}^{C}_{uv}\stackrel{{\scriptstyle\mbox{def.}}}{{=}}\,A_{uv}\mathcal{L}^{*}_{C}\Big{(}\sum_{t=1}^{T}\,\operatorname{SteepSigmoid}(\mathcal{T}(\hat{x}_{u})_{t}-\mathcal{T}(\hat{x}_{v})_{t})\Big{)},$ (9)
@@
where $\mathcal{L}^{*}_{C}$ is a function that takes as input the expression above and $\operatorname{SteepSigmoid}(x)=\frac{1}{1+e^{-10x}}$ (we omit some details here for simplicity, see Appendix C.3 for details). The loss for two causally connected events $u\preccurlyeq v$ in the first neighborhood of each other ($A_{uv}=1$) is minimized when $\hat{x}_{u}\lesssim^{\mathcal{T}}\hat{x}_{v}$ is satisfied (equation 5). If the time coordinates of $\hat{x}_{v}$ associated with the event $v$ are all greater than those of $\hat{x}_{u}$ for $u$, then all $\operatorname{SteepSigmoid}$ activation functions will return $\approx 0$.
Table 1: DAG embedding results. Embedding dimension $D=T=2,4,10$.
Metric Embedding Dim Distortion (average $\pm$ std) Max Distortion Directionality
$\|\mathbf{x}-\mathbf{y}\|^{0.5}\log(1+\|\mathbf{x}-\mathbf{y}\|)^{0.5}$ 2 1.09 $\pm$ 0.24 3.18 1.0 Neural Spacetime
4 1.02 $\pm$ 0.06 1.51 1.0
10 1.00 $\pm$ 0.03 1.24 1.0
4 1.02 $\pm$ 0.07 1.75 1.0
10 1.00 $\pm$ 0.04 1.47 1.0
4 1.11 $\pm$ 0.41 8.92 1.0
10 1.01 $\pm$ 0.05 1.31 1.0
$\|\mathbf{x}-\mathbf{y}\|^{0.5}\log(1+\|\mathbf{x}-\mathbf{y}\|)^{0.5}$ 2 2.86 $\pm$ 5.22 72.66 0.99 Minkowski
4 1.70 $\pm$ 2.77 71.09 0.99
10 1.21 $\pm$ 1.33 35.58 0.99
4 1.70 $\pm$ 5.21 77.03 0.99
10 1.19 $\pm$ 1.09 25.18 0.99
4 2.49 $\pm$ 8.72 198.04 0.98
10 1.18 $\pm$ 2.49 82.67 0.99
$\|\mathbf{x}-\mathbf{y}\|^{0.5}\log(1+\|\mathbf{x}-\mathbf{y}\|)^{0.5}$ 2 $\infty$ $\pm$ $\infty$ 0.99 De Sitter
4 -4.33 $\pm$ 816.47 10235.71 0.99
10 288.17 $\pm$ 9794.97 324027.5 0.99
4 174.69 $\pm$ 3637.32 115851.88 0.99
10 -10.66 $\pm$ 739.71 8997.62 0.99
4 83.04 $\pm$ 4313.82 97524.21 0.94
10 418.26 $\pm$ 6599.80 150543.73 0.94
4 Experimental Results
@@
Synthetic Weighted DAG Embedding. We generate DAGs (see Appendix D.2) embedded in the 2D plane with associated local distances between nodes given by several metrics. We measure the embedding capabilities of NSTs compared to closed-form spacetimes such as the Minkowski and De Sitter spaces in Law & Lucas (2023). Although we do not optimize the geometries in the case of the baselines, we do use a neural network encoder to map points to events in the manifolds. We quantify both the average and maximum metric distortion (the ratio between true and predicted distances) following Kratsios et al. (2023b), as well as the accuracy of the time embedding. As can be seen in Table 1, we are always able to embed edge directionality (0 for no edges embedded correctly, 1 for all edges embedded correctly). In terms of metric distortion, as the embedding dimension increases, both average and maximum distortion decrease. NSTs are particularly good at retraining low distortions in low-dimensional embedding spaces. See Table 9 for more results.
Table 2: Embedding results for real-world web page hyperlink and gene regulatory networks.
Neural Spacetime Minkowski De Sitter
Dataset
Embed. Dim
Distortion
Max Distortion
Directionality
Distortion
Max Distortion
Directionality
Distortion
Max Distortion
Directionality
Cornell 2 1.00 $\pm$ 0.07 1.31 0.93 1.07 $\pm$ 0.70 9.43 0.94 -55.83 $\pm$ 890.45 3950.88 0.92
4 1.00 $\pm$ 0.04 1.08 0.94 1.00 $\pm$ 0.00 1.01 0.94 -20.60 $\pm$ 249.49 403.46 0.94
10 1.00 $\pm$ 0.04 1.08 0.94 1.00 $\pm$ 0.00 1.00 0.94 0.80 $\pm$ 126.26 1543.07 0.93
Texas 2 1.01 $\pm$ 0.10 2.27 0.89 1.12 $\pm$ 1.73 31.27 0.90 -0.29 $\pm$ 84.42 818.10 0.90
4 1.00 $\pm$ 0.01 1.05 0.90 1.00 $\pm$ 0.00 1.00 0.90 42.03 $\pm$ 795.51 13939.25 0.90
10 1.00 $\pm$ 0.00 1.00 0.90 1.01 $\pm$ 0.01 1.05 0.90 2.60 $\pm$ 70.33 1107.60 0.90
Wisconsin 2 1.00 $\pm$ 0.10 1.67 0.89 5.07 $\pm$ 65.99 1410.03 0.90 2.06 $\pm$ 63.46 1291.31 0.89
4 1.00 $\pm$ 0.04 1.16 0.89 1.00 $\pm$ 0.04 1.19 0.90 -0.78 $\pm$ 27.91 114.24 0.90
10 1.00 $\pm$ 0.04 1.20 0.89 1.13 $\pm$ 0.69 16.28 0.90 0.04 $\pm$ 215.94 2862.19 0.89
In silico 2 1.06 $\pm$ 0.47 18.54 1.00 105.42 $\pm$ 4671.85 209248.72 0.94 -63.59 $\pm$ 1866.69 56626.97 0.92
4 1.00 $\pm$ 0.09 1.73 1.00 0.25 $\pm$ 54.57 1315.76 0.95 -468.81 $\pm$ 33021.14 65289.22 0.92
10 1.00 $\pm$ 0.05 1.32 1.00 1.00 $\pm$ 0.05 3.69 0.99 -129.13 $\pm$ 9623.30 261531.59 0.93
E. coli 2 1.02 $\pm$ 0.45 15.37 1.00 -4.25 $\pm$ 149.61 438.34 0.97 34.65 $\pm$ 2637.50 119047.23 0.91
4 1.00 $\pm$ 0.06 2.62 1.00 1.00 $\pm$ 0.01 1.08 0.98 -2.00 $\pm$ 3294.81 130509.59 0.91
10 1.00 $\pm$ 0.05 1.17 1.00 1.00 $\pm$ 0.01 1.01 0.99 -8.26 $\pm$ 94.57 652.96 0.91
S. cerevisiae 2 1.05 $\pm$ 0.34 10.18 1.00 -2.38 $\pm$ 173.57 151.43 0.91 55.36 $\pm$ 3960.09 160278.39 0.90
4 1.00 $\pm$ 0.07 1.63 1.00 1.04 $\pm$ 2.25 63.17 0.98 -28.60 $\pm$ 1175.67 63086.54 0.90
10 1.00 $\pm$ 0.05 1.57 1.00 1.01 $\pm$ 0.02 1.39 0.99 -121.17 $\pm$ 7550.16 84724.25 0.91
@@
Real-World Network Embedding. We test our approach on real-world networks. In Table 2, we present results for the Cornell, Texas, and Wisconsin (WebKB) datasets (Rozemberczki et al., 2021), which are based on webpages represented as nodes and directed hyperlinks between them. All the nodes have bag-of-word features that we use as input of the neural spacetime encoder. The neural partial order encodes the hyperlink directionality between websites, and the neural (quasi-)metric learns the connectivity strength as the cosine similarity between connected webpage features. We achieve very low distortions, showcasing the embedding capabilities of our network. Note that, from a metric learning perspective, real-world datasets are generally less challenging than the metrics presented in Table 1, which we chose to be particularly unconventional on purpose. In terms of the time embedding, we manage to mostly encode directionality; however, these datasets are not pure DAGs since they contain some directed cycles, so it is not possible to embed them perfectly. We also work with real-world gene regulatory network datasets (Marbach et al., 2012) in line with spacetime representation learning literature (Law & Lucas, 2023; Sim et al., 2021). We achieve good spatial and causal embeddings for the In silico, Escherichia coli, and Saccharomyces cerevisiae datasets, see Table 2. Experimental details and hyperparameters for all setups can be found in Appendix D. Additionally, we present tree (spatial only) embedding experiments comparing neural snowflakes to neural (quasi-)metrics in NSTs, as well as hyperbolic neural networks (HNNs), in Appendix D.1.